If a and b are real numbers, then the product of a+bi and a-bi is
(1) always imaginary
(2) always a real number
(3) sometimes,but not always, a real number
(4) sometimes,but not always, imaginary
----------------------------------------------------------------------------------
Please help choose the correct answer choice and please explain how to get the answer Thank you

Answer by stanbon(48508)

If a and b are real numbers, then the product of a+bi and a-bi is
(1) always imaginary
(2) always a real number
(3) sometimes,but not always, a real number
(4) sometimes,but not always, imaginary
----------------
(a+bi)(a-bi) = a^2 +abi -abi -(bi)^2
= a^2 + b^2
and this is always a Real Number because
a and b are always Real Numbers.
=========================================
Cheers,
Stan H.